OpenStudy (anonymous):

How in the world do I solve 6^4√2-14^3√2+5^3√2+7^4√2 :( my answer choices are a) 13^4√-9^3√4 b) 13^4√2-9^3√2 c) 4√2 d) 4√4

OpenStudy (anonymous):

Woah

OpenStudy (anonymous):

Yea... Algebra 2 online is aweful

OpenStudy (anonymous):

try using pemdas, just work through it i guess

OpenStudy (amistre64):

you can start by changing all the sqrt(2) parts to something more readable like a or b or something

OpenStudy (amistre64):

are those sqrts in the exponent?

OpenStudy (anonymous):

That doesnt make an sense :( I'm so confusedd !! the exponent is infront of the square root sign

OpenStudy (amistre64):

\[ 6^{4a}-14^{3a}+5^{3a}+7^{4a}\]

OpenStudy (amistre64):

or \[ 6^4a-14^3a+5^3a+7^4a\]???

OpenStudy (anonymous):

\[6^{4}\sqrt{2}-14^{3}\sqrt{2}+5^{3}\sqrt{2}+7^{4}\sqrt{2}\]

OpenStudy (amistre64):

ahh, then since i cleaned it up with as, you can factor it out and you can modify the base exponent stuff with logs most likely

OpenStudy (anonymous):

logs...?

OpenStudy (amistre64):

\[6^4a-14^3a+5^3a+7^4a\] \[a(6^4-14^3+5^3+7^4)\] getting all this in like base terms is useful, and thats what logs can do maybe

OpenStudy (amistre64):

or, expand them all out and add em up

OpenStudy (amistre64):

since i got no idea what level of material and study you are at, im pretty much looking at this thru my own concepts

OpenStudy (anonymous):

Algebra 2...

OpenStudy (anonymous):

Would my answer be letter B?

OpenStudy (amistre64):

6^4a-14^3a+5^3a+7^4a = 13^4a-9^3a lets give it a shot http://www.wolframalpha.com/input/?i=sqrt%282%29%286%5E4-14%5E3%2B5%5E3%2B7%5E4+-13%5E4%2B9%5E3%29 im gonna say no

OpenStudy (anonymous):

D: but nothing is matching my answer choices!!

OpenStudy (anonymous):

I'm just going to guess on this one :(

OpenStudy (amistre64):

6^4−14^3+5^3+7^4 = 1028 which option equals 1028 if you ignore the sqrt